OpenStudy (anonymous):

How many sides does a regular polygon have if each exterior angle measures 20

OpenStudy (anonymous):

i hope this will help. . http://answers.yahoo.com/question/index?qid=20071022065937AAYgfNy

OpenStudy (anonymous):

360/20=18

OpenStudy (anonymous):

20 = 360/x 20x = 360 x = 360/20 = 18